OpenStudy (anonymous):

True or false? When you simplify each side of the equation 4(x + 3) = 5 – (x + 1), you will get the equivalent equation 4x + 12 = -x + 4.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

truee

OpenStudy (anonymous):

-(x+1) = -x-1

OpenStudy (anonymous):

basics

OpenStudy (anonymous):

4x+12=-5x+5